OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 512 cm3?

OpenStudy (anonymous):

i dunno mang, maybe try cube rooting it. dam this is hard

OpenStudy (anonymous):

\[\sqrt[3]{512} = 16 \sqrt{2}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Um...\[\sqrt[3]{512}=8=>8^3=512\]

OpenStudy (anonymous):

I used my calculator and got 16 * sqrt(2)

OpenStudy (anonymous):

I did not get 8 as a result. Is \[16\sqrt{2} = 8\]

OpenStudy (anonymous):

Well try cubing 16sqrt2, do you get 512? I got 11585.2375 8^3=512 and when cube root 512 you get 8. 8*8=64 =>64*8=512 thus 8*8*8=512

OpenStudy (anonymous):

Oops I made a mistake. 512 is not 16sqrt(2) cubed but everybody makes mistakes sometimes.

OpenStudy (anonymous):

No problem, just wanted to point it out